What is the net force on a A 1400 kg car speeds up from 12.3 m/s to 37.3 m/s in 8.34 s. What was the net force acting on the car?



1
Expert's answer
2021-01-13T11:37:41-0500

Let's first find the acceleration of the car:


"a=\dfrac{v-v_0}{t}=\dfrac{37.3\ \dfrac{m}{s}-12.3\ \dfrac{m}{s}}{8.34\ s}=3.0\ \dfrac{m}{s^2}."

Finally, we can find the net force on a car from the Newton's Second Law of Motion:


"F_{net}=ma=1400\ kg\cdot 3.0\ \dfrac{m}{s^2}=4200\ N."

Answer:

"F_{net}=4200\ N."